Please help me find the answe tho these questions: A triangle ABC with A(1, 1) and B(-1,4). The gradients of AB, AV and BC are -3m, 3m and m respectively.
a) Find the value of m
b) Find the coordinates of C
c) Show that AC = 2AB [10 Marks]

Answer by math_tutor2020(3817)


Part a)

I'm assuming you meant to say AC instead of AV.

A = (1,1)
B = (-1,4)
Gradient = slope
Slope = (change in y)/(change in x)
Slope = (y2-y1)/(x2-x1)
Slope of AB = (4-1)/(-1-1)
Slope of AB = (3)/(-2)
Slope of AB = -3/2

Set this equal to the stated slope of AB (which was -3m) and solve for m
-3/2 = -3m
m = (-3/2)*(-1/3)
m = 1/2

---------
Part b)

Let C be located at (x,y).
The goal is to find the numbers that replace x and y.

A = (1,1)
C = (x,y)
Slope of AC = (y2-y1)/(x2-x1)
Slope of AC = (y-1)/(x-1)
Set this equal to 3m. Then plug in m = 1/2

Slope of AC = 3m
(y-1)/(x-1) = 3m
(y-1)/(x-1) = 3(1/2)
(y-1)/(x-1) = 3/2

Let's now cross multiply and isolate y
(y-1)/(x-1) = 3/2
2(y-1) = 3(x-1)
2y-2 = 3x-3
2y = 3x-3+2
2y = 3x-1
y = (3x-1)/2
y = (3/2)x - (1/2)
y = 1.5x - 0.5
This is the equation of line AC.

Next, we'll find the slope of line BC
B = (-1,4)
C = (x,y)
Slope of BC = (y2-y1)/(x2-x1)
Slope of BC = (y-4)/(x-(-1))
Slope of BC = (y-4)/(x+1)

Set this equal to the stated slope of m
m = (y-4)/(x+1)

Then we'll plug in these items
m = 1/2 = 0.5
y = 1.5x - 0.5

We end up with this
m = (y-4)/(x+1)
0.5 = (1.5x-0.5-4)/(x+1)
0.5 = (1.5x-4.5)/(x+1)

It's a lot of steps to take in (there might be a faster more efficient way, so I'll let another tutor step in for that).
However, it shouldn't be too bad once you get enough practice with algebra.

From here, we isolate x to get...
0.5 = (1.5x-4.5)/(x+1)
0.5(x+1) = 1.5x-4.5
0.5x+0.5 = 1.5x-4.5
0.5+4.5 = 1.5x-0.5x
5 = x
x = 5
Then use this x value to find its paired y value
y = 1.5x-0.5
y = 1.5*5-0.5
y = 7.5-0.5
y = 7
Point C is located at (x,y) = (5, 7)
Don't forget to use parenthesis when writing the coordinates.

We have these three point locations
A = (1,1)
B = (-1,4)
C = (5,7)
I'll let you use the slope formula to confirm that -3m, 3m and m represent the slopes of AB, AC, and BC respectively (where m = 1/2 = 0.5 calculated earlier).
A tool like GeoGebra is very handy to quickly confirm we have the correct values so far.

---------
Part c)

Focus on points A and C. Use the distance formula to find the length of the segment AC.
A = (x1,y1) = (1,1)
C = (x2,y2) = (5,7)


















Segment AC is exactly units long.
If we can show that segment AB is exactly units long, then we'd be able to wrap things up.

Use the distance formula for points A and B
A = (x1,y1) = (1,1)
B = (x2,y2) = (-1,4)














We see that AB is exactly units long, so the proof is concluded.
We have shown that AC is twice as long as AB.

===============================================================================

Answers:
a) m = 1/2
b) Point C is at (5, 7)
c) Proof is shown above